Prev Next

SW&IT Meeting Minutes 2018-04-23

Attendees: Andrew, David, Doug, Igor, Joe, Kellee, Matthieu, Petr, Stephen, Xianyu, Yang

Previous Action Items

  • NetBrain (David - schedule demo)
  • VM/SSD (David - followup meeting)
  • Scriptor update webpage message (Stephen)
  • Snapshot Argos tool (Matthieu) missing doc, howto, experts only ? Tomaso who is leaving ? Doug Miller

News from management

  • board meeting
    • status of instruments, commissioning and handovers (LUCI, MODS, ARGOS)
    • future plans LBTI, support 2018B
    • iLocater AI from PDR
    • PEPSI probe interest for facility
    • SOUL, lean LN, 2ng Gen instruments update and plans
    • Scientific return
    • TOO policy
    • SAC -> STAC
    • new user committee
    • promote Q
    • Management scheme (Program manager)
    • RCSA transition
    • Safety committee
    • Budget, etc.

  • Labor Record Keeping WG
    provide valuable data for
    (a) conveying actual labor expenditures - time and/or cost- to stakeholders,
    (b) comparing actual to estimated labor expenditures,
    (c) improving forecasting of future labor expenditures, and
    (d) assessing staffing requirements

  • OC
    • board meeting debrief
    • Water mitigation
    • SSD planning, first up time phasing PB/AO
    • Meetings structure and duplication (IT, WPC, Inst, etc.)
    • Argos SO remote support, progress

  • team calendar, sick time

Outlook

Team reports

AC

Last week
  • Created exception for "lbtElevationDeg is None" for LAT and forward it to arbitrator and LAT controller GUIs
  • Depolyed observer status GUI
  • Added increment and decement loop feature to spiral search
  • Removed "Optimize Gain" button on the arbitrator and added "Tip Tilt Only". Sets and applies gains TT=1, LO=0, HI=0.
  • Fixed Vibration Analyzer tool
  • Closed IT#7131
This week
  • ARGOS run - add plots to help determine optimal focus range for sharpen function

DE

  • Last week
    • Set up Veeam bulk backup and notifications
    • Requested quote for Veeam (4x socket license)
    • Discovered some VMWare licensing anomalies; should not be a problem for our use case
    • Set up a clone of the PIT VM on the mountain to work around planned outages this week
  • This week
    • Follow up on Veeam licence quote
    • Get replacement for Laura's laptop--too many problems to be acceptable
    • Test Veeam restore
    • Any last-minute items to cover for loss of network access to the mountain?

DS

Last Week:
  • ARGOS pointing models & backtests in prep for upcoming run
  • Some support to Matthieu on mgt transition tasks
  • DIMM calibration automation research
This Week
  • Continued work on understanding why pointing is so bad - searching for systematics
  • Possibly some DIMM focus automation design thoughts

IS

Last Week:
  • Working on a current view for the LUCI scheduler. I've talked to Dave T. about design and functionality. About 60% done, ETA ~2days
  • Looking at how LUCI software deals with mask names and slot numbers. Send email to Alex and got some insights but no luck yet.
  • The Italian partner wants a PIT for this proposal season but we need to special case for them. We'll try to release Italian PIT ASAP by early next week.
  • OT Scripting: David E. made me a clone of the pit.lbto.org to be run on the mountain. OT scripts in the obs# machines will use this VM, local to the mountain, instead of the one in Tucson. This is achieved by adding an entry to the /etc/hosts file in the obs# machines that points the DNS otscripting.lbto.org to the mountain IP address. This is ahead of the two day intermittent network outage to happen early next week.

This Week:
  • Italian PIT
  • LUCI scheduler
  • LUCI masks if time allows
  • Vacation on Friday

JA

Last week
  • Planner Synchronous Offsets
    • Meetings with ME, Dave Thompson - preliminary test cases
    • Implemented to a first order (they group correctly but need further testing)
This week
  • Planner -
    • Overheads for LBC, MODS
    • "Business logic" validation
    • Testing of synchronous offsets (Will probably need to meet with ME to think up more use cases)
    • Testing of functionality against original specification
  • Need test release for OT mask testing...

KS

Last week
  • Looked at couple of DIMM core files from the weekend (shutdown race conditions)
  • Started documentation updates for OAC HK upgrade
  • Figured out the argument issues with the OAC client programs
  • SVN issues - are AO INDI drivers in?
    - what's in tools repository?
  • Looked at GCS logging IT - trouble getting TCS configuration correct on the remote server here, but it's ok now

This week
  • GCS patch for logging issue
  • Review DIMM NDA and other 3rd party issues in my SVN repositories
  • OAC HK upgrade:
    • handle errors from EPICS writes
    • add the harder alarming cases - if it's off, but wasn't commanded by OAC to be off
    • make a "poweronly" option for the getdata 8051 board

MB

  • management OC, WPC
  • met Argos, iLocater
  • mountain trip:
    • attended OC on arrival (!)
    • saw laser (on gui and webcam)
    • learnt a lot about Argos, TAN (twice as nice)
    • more discussions with Jose Borelli (inc. DIMM, irc, some history too)
    • sneak peek OSA, telescope, control room, mountain and night-time activities in general
    • met Leroy and Chris eventually on Friday, got a tour and better understanding of the telescope
  • repos browsing
  • git migration kickoff
  • batch triage 250+ safety ITs for DC
  • create new osa- scripts repos

Next:
  • 'activate' first migrations
  • repo browsing, tcsweb, tools, etc.
  • responsibilities and assignments draft

PK

  • IT 7190 - LSS backlog, probably tracked to TIME_WAIT sockets, worked on graphing various plots to confirm that
  • fixed IT 7218 (co-pointing limit when NS tracking) for DETXY, discover the same applies to Ra/Dec offsets (not yet fixed due to different nature of those), discover race conditions in PCS code when Ephemeris update is enabled

Next:
  • continue on IT 7218 - fix race condition, prepare patch for installation, get permission to install (PCS can crash mid-time while performing asynchronous offsets for Quasy, so better to fix)
  • continue on IT 7190, get the proofs
  • Time to Limit (IT 3xxx)

SH

Unfinished from last week:
  • cisrv01 job configuration bootstrap and git integration * git sparse checkouts
  • projector lamp in 505
  • finish moving developer VMs to SSD
This week:
  • wiki.lbto.org -- add the mountain
    • TWiki RPM and kickstart
    • DNS changes
    • Just read-only up there for now?
  • cissrv01 to build system's RPMs
  • camera page problems reported by Gustavo.

XZ

Late week:
  • Meeting OVMS+ on ASM, powerbackplane status, MOXA issues, and AO meeting
  • FLAO/ ARGOS TT LUT support
  • LUCI ARGOS auto data collection routines debug and test
  • Found a solution for IT7213 (Connection to RTDB lost, AdSec Arbitrator unresponsive)
  • Created and put error information for IT7221 (SX stageZ home issue)
  • Fixed the ice issue between AOS UAO by simulators
This week:
  • Fix IT7213
  • Continue working on the power backplane
  • Technical discussion with Roberto (MG) and Guido?

YZ

  • Compiled the AO code on the AO VM using GCC version 5, 6, and 7. All seem to be working (with minor code changes, which still works under GCC 4.4). This needs to be further verified. So far it seems promising. If indeed successful, this opens up a lot more opportunities in using C++11 and beyond features.
  • Modified the "mergelogs" tool to work with current Boost library interface. Successfully recompiled it on CentOS 7 with default GCC and default Boost installation. Although I didn't try to fix the bugs in "mergelogs". Current obvious bugs seem to be related to crashes with file processing involving multiple threads.
  • Developing prototype logging facilities that enables a ordered log message output from multiple threads (loggers) to a single file output. Currently implemented as a component in the "spdlog" library. Further tuning maybe needed. If prototype is deemed okay, then will redevelop the AO logging facility based on this.
  • A brief day trip to the LBT site to visit the telescope and the building.

Action items

  • xxx
  • xxx
  • xxx

-- %USERSIG{MatthieuBec}%

Comments

-- %BUBBLESIG{XianyuZhang - 2018-04-23}%
 
Topic revision: r26 - 10 Dec 2018, UnknownUser
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback